### v1

# RSCaMa: Remote Sensing Image Change Captioning with State Space Model

Abstract
Remote Sensing Image Change Captioning (RSICC) aims to identify surface changes in multi-temporal remote sensing images and describe them in natural language. Current methods typically rely on an encoder-decoder architecture and focus on designing a sophisticated neck to process bi-temporal features extracted by the backbone. Recently, State Space Models (SSMs), especially Mamba, have demonstrated outstanding performance in many fields, owing to their efficient feature-selective modelling capability. However, their potential in the RSICC task remains unexplored. In this paper, we introduce Mamba into RSICC and propose a novel approach called RSCaMa (Remote Sensing Change Captioning Mamba). Specifically, we utilize Siamese backbones to extract bi-temporal features, which are then processed through multiple CaMa layers consisting of Spatial Difference-guided SSM (SD-SSM) and Temporal Traveling SSM (TT-SSM). SD-SSM uses differential features to enhance change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner. Experimental results validate the effectiveness of CaMa layers and demonstrate the superior performance of RSCaMa, as well as the potential of Mamba in the RSICC task. Additionally, we systematically compare the effects of three language decoders, including Mamba, GPT-style decoder with causal attention mechanism, and Transformer decoder with cross-attention mechanism. This provides valuable insights for future RSICC research. The code will be available at https://github.com/Chen-Yang-Liu/RSCaMa
Index Terms:
Change captioning, Mamba, State Space Model, Spatial Difference-guided SSM, Temporal Traveling SSM.
I Introduction
Multi-temporal remote sensing images captured by satellites provide rich data resources for surface dynamic monitoring. Remote Sensing Image Change Captioning (RSICC) uses multi-temporal images of the same area to identify and describe surface changes in the form of natural language, such as the categories and attributes of changing objects and the positional relationships between objects. It greatly enhances the readability and practicality of change interpretation information, and has significant application in various fields, including land planning [1], disaster detection [2], and urban expansion research [3].
RSICC involves remote sensing image processing and natural language generation. Current mainstream methods typically employ an encoder-decoder structure. In the encoder, a convolutional neural network (CNN) or Vision Transformer (ViT) is used as the backbone to extract visual features from bi-temporal images, a well-designed “neck” to strengthen the spatial-temporal correlation of bi-temporal features and capture interesting changes features. Finally, the decoder converts these visual features into captions by using language models such as recurrent neural networks or the Transformer.
Most current research focuses on improving the design of the neck module, usually using CNN or Transformer as the basic component, and integrating attention mechanisms to improve the accuracy of change captioning. Hoxha et al. [4] firstly explored the RSICC task and proposed two CNN-based bi-temporal fusion strategies. Liu et al. [5] constructed a large dataset LEVIR-CC and benchmarked several methods such as DUDA [6], MCCFormer [7]. Besides, they proposed a dual-branch Transformer to locate change areas and a Multistage Bitemporal Fusion (MBF) module to perform the fusion of multi-scale features. Chang et al. [8] proposed a hierarchical self-attention module to locate change-related features. PSNet [9] leverages scale-aware reinforcement modules to enhance the perception of changed objects of varying sizes. PromptCC [10] uses an image-level classifier and a feature-level encoder to decouple the RSICC task. Besides, a multi-prompt learning strategy is proposed to exploit a large language model for captioning effectively. CT-Net [11] use employs a Multi-Layer Adaptive Fusion (MAF) module to minimize the impact of irrelevant visual features and a Cross Gated-Attention (CGA) module to attend to multi-scale features during the caption-generation process.
Recently, State Space Models (SSMs) [12] with linear complexity, especially Mamba [13], have demonstrated excellent performance in many language and vision tasks due to their efficient contextual modelling capabilities. In the field of remote sensing, some research has explored the application of SSMs. For instance, Pan-Mamba [14] for pan-sharpening, SpectralMamba [15] for Hyperspectral image classification, MiM-ISTD [16] for infrared small target detection, Samba [17] and RS-Mamba [18] for image segmentation. Two surveys of SSMs can be found in [19, 20].
However, the potential of SSM, particularly Mamba, remains underexplored in the RSICC task. In this paper, we introduce Mamba into the RSICC and propose RSCaMa (Remote Sensing Change Captioning Mamba). Specifically, we utilize the Siamese backbones to extract bi-temporal features, which are subsequently fed into multiple CaMa layers. In each CaMa layer, we propose Spatial Difference-guided SSM (SD-SSM) to enhance spatial change perception, and Temporal Traveling SSM (TT-SSM) to perform temporal interaction between bi-temporal features. The bi-temporal features processed by the CaMa layers will be passed through a projection layer to aggregate bi-temporal information and narrow the gap between the visual and language domains. Finally, language decoders generate sentences describing changes of interest. We experimentally confirm the effectiveness of RSCaMa and demonstrate the potential of Mamba as a new model design component in the RSICC task. Additionally, we compare the effect of three language decoders: Mamba language model, GPT-style decoder with causal attention, and Transformer decoder with cross-attention, providing valuable insight for future RSICC research.
Our contributions can be summarized as follows:
  • We introduce SSM into the RSICC, a multi-modal task, and propose RCaMa for efficient spatial-temporal modelling, providing a benchmark for Mamba-based RSICC.
  • We propose the CaMa layer consisting of SD-SSM and TT-SSM. SD-SSM uses differential features to enhance change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner.
  • The experiment shows the RSCaMa’s SOTA performance and the potential of Mamba in the RSICC task. Besides, we systematically evaluate different language decoding schemes, providing valuable insight for future research.
Figure 1: The illustration of the proposed RSCaMa, which consists of three main components: the backbone, multiple CaMa layers, and the language decoder. The CaMa layers play a pivotal role in significantly enhancing the extraction of discriminative features by integrating SD-SSM and TT-SSM. Specifically, SD-SSM focuses on enhancing change perception in spatial features, while TT-SSM concentrates on temporal interaction.
II Methodology
In this section, we present an in-depth explanation of our proposed RSCaMa. We commence by introducing the foundation of the state space model. Following that, we offer a comprehensive overview of the RSCaMa model and delineate its principal constituents: SD-SSM and TT-SSM.
II-A Preliminaries
SSMs  [12] represent a class of sequence models with linear complexity, inspired by linear time-invariant (LTI) continuous system mapping a one-dimensional function or sequence $x(t)\in\mathbb{R}$ to $y(t)\in\mathbb{R}$ through an implicit hidden state $h(t)\in\mathbb{R}^{N}$. The system can be mathematically expressed as follows:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)$ (1)
$y(t)=\mathbf{C}h(t)$ (2)
where $h^{\prime}(t)$ refers to the time derivative of $h(t)$, ${A}\in\mathbb{R}^{N\times N}$ is the evolution parameter, ${B}\in\mathbb{R}^{N\times 1}$ and ${C}\in\mathbb{R}^{1\times N}$ serve as the projection parameters.
To process discrete sequences and map the input sequence $\{x_{1},x_{2},...,x_{K}\}$ to the output sequence $\{y_{1},y_{2},...,y_{K}\}$, SSMs leverage zero-order hold (ZOH) discretization to convert continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into their discrete counterparts $\bar{\mathbf{A}}$ and $\bar{\mathbf{B}}$. The transformation is achieved by incorporating a timescale parameter $\mathbf{\Delta}\in\mathbb{R}^{D}$ and can be formally defined as follows:
$\bar{\mathbf{A}}$ $=\mathbf{exp(\Delta A)}$ (3)
$\bar{\mathbf{B}}$ $=\mathbf{(\Delta A)}^{-1}\mathbf{(exp(\mathbf{\Delta}A)-I)\cdot\mathbf{\Delta}B}\approx\mathrm{\mathbf{\Delta}}B$ (4)
Then, the discrete representation of the linear system can be formulated as follows:
$h_{k}=\bar{\mathbf{A}}h_{k-1}+\bar{\mathbf{B}}x_{k}$ (5)
$y_{k}=\mathbf{C}h_{k}$ (6)
Mamba [13] breaks the LTI limitation of SSM by introducing a selective state space mechanism. It allows the matrices $\bar{\mathbf{B}}$, $\bar{\mathbf{C}}$, and $\mathbf{\Delta}$ to vary with the input sequences, enhancing selective information processing across sequences, enabling the model to be contextually aware of the input.
II-B RSCaMa
As depicted in Fig. 1, we present the overall architecture of the proposed RSCaMa. RSCaMa consists of three main components: the backbone, multiple CaMa layers, and the language decoder. Following prior research [10], we employ the CLIP image encoder that has been image-text aligned as the backbone to extract bi-temporal features, which are then added with position embeddings and tokenized as $I_{T_{i}}\in\mathbb{R}^{L\times D}(i=1,2)$. $L$ and $D$ are the number and dimensions of tokens respectively. Subsequently, they are fed into multiple CaMa layers. The CaMa layers leverage the synergistic effect of SD-SSM and TT-SSM to further enhance discriminative feature extraction, with SD-SSM enhancing spatial change perception and TT-SSM concentrating on temporal interaction. The two token sequences are iteratively processed through multiple CaMa layers and then input into a projection layer to further bridge the gap between the visual and language domains. Finally, the language decoder utilizes the resulting embeddings to generate descriptive sentences. For the RSCaMa with $N$ CaMa layers, assuming that the bi-temporal features from the backbone are $I_{T_{1}}^{(0)}$ and $I_{T_{2}}^{(0)}$, the flow of caption generation is expressed as follows:
${I_{T_{1}}^{(l)},I_{T_{2}}^{(l)}}$ $=\mathrm{\Phi_{CaMa}^{(l)}}(I_{T_{1}}^{(l-1)},I_{T_{2}}^{(l-1)}),l=(1,2,...N)$ (7)
$I_{F}$ $=\mathrm{\Phi_{Proj}}([I_{T_{1}}^{(N)};I_{T_{2}}^{(N)}])$ (8)
$\mathrm{cap}$ $=\mathrm{\Phi_{Dec}}(I_{F})$ (9)
where [;] denotes feature concatenation operation on the channel dimension. $\mathrm{\Phi_{Proj}}(\cdot)$ represents a projection layer consisting of a linear layer and a residual convolution block.
Figure 2: The structure of the SD-SSM. We multiply the differencing features and the output of bidirectional SSMs to improve the change-awareness.
II-C SD-SSM
Fig. 2 presents the structure of our proposed SD-SSM. The standard Mamba is designed for one-dimensional sequences. Its causal scanning mechanism makes it unsuitable for spatial-aware understanding in visual tasks. Inspired by [21], we perform Forward-SSM and Backward-SSM scans on flattened visual tokens to achieve bidirectional modelling, enhancing spatial understanding of visual features. Furthermore, to improve the model’s perception of changes, we multiply the bi-temporal differencing features and the output of bidirectional SSMs to guide the model. Let’s denote bi-temporal features input to SD-SSM as $I_{T_{1}}$ and $I_{T_{2}}$, the processing of SD-SSM can be represented as follows:
$I^{o}_{T_{i}}$ $=\mathrm{\Phi_{L}}(I^{F}_{T_{i}})+I_{T_{i}},i=1,2$ (10)
$I^{F}_{T_{i}}$ $=I^{f}_{T_{i}}*I_{d}+I^{b}_{T_{i}}*I_{d}$ (11)
$I_{d}$ $=\sigma(\mathrm{\Phi_{L}}(I_{T_{2}}-I_{T_{1}}))$ (12)
$I^{f}_{T_{i}}$ $=\mathrm{\Phi_{SSM}}(I^{p}_{T_{i}})$ (13)
$I^{b}_{T_{i}}$ $=\mathrm{flip}(\mathrm{\Phi_{SSM}}(\mathrm{flip}(I^{p}_{T_{i}})))$ (14)
$I^{p}_{T_{i}}$ $=\sigma(\mathrm{\Phi_{Dwc}(\Phi_{L}(\Phi_{Norm}}(I_{T_{i}}))))$ (15)
where $I^{o}_{T_{i}}$ is the output, $\sigma$ is the SiLU activation function, $\mathrm{\Phi_{Dwc}}(\cdot),\mathrm{\Phi_{L}}(\cdot),\mathrm{\Phi_{Norm}}(\cdot)$ respectively represent depth-wise convolution, linear projection, and normalization. $\mathrm{flip}(\cdot)$ represents reversing the order of a sequence.
Figure 3: The structure of the TT-SSM, which rearranges two sequences in a bi-temporal token-wise interleaving manner to facilitate temporal interaction.
II-D TT-SSM
SD-SSM spatially interacts with bi-temporal features and differencing features to promote change perception. Furthermore, we propose TT-SSM to perform temporal interaction to further enhance the bi-temporal features. Fig. 3 shows the structure of the TT-SSM. TT-SSM achieves bi-temporal feature interaction by scanning the two token sequences in the time dimension. Specifically, for the two sequences $I_{T_{1}}\in\mathbb{R}^{L\times D}$ and $I_{T_{2}}\in\mathbb{R}^{L\times D}$ from SD-SSM, we rearrange them in a bi-temporal token-wise interleaving manner to construct sequences $I_{s}=[I_{T_{1}}^{(0)},I_{T_{2}}^{(0)},I_{T_{1}}^{(1)},I_{T_{2}}^{(1)},I_{T_{1}}^{(2)},...]\in\mathbb{R}^{2L\times D}$ that effectively facilitate temporal information exchange. Finally, $I_{s}$ will be processed by bidirectional SSM and subsequently split into bi-temporal sequences in the original order. After that, the two sequences will be sent to the next CaMa layer and processed iteratively by SD-SSM and TT-SSM.
III Experiments
III-A Experimental Setting
III-A1 Dataset
LEVIR-CC [5] is a large-scale RSICC dataset comprising 10,077 pairs of images. These images capture changes over time in 20 distinct regions across Texas, USA, with a temporal range spanning from 5 to 15 years. Each image pair of the dataset has five annotated captions, totaling 50,385 descriptive captions, describing diverse ground surface transformations like building, road, and vegetation.
III-A2 Evaluation Metrics
We employ four quantitative metrics to measure the captioning accuracy, including BLEU-N (N=1,2,3,4), ROUGEL, METEOR, and CIDEr-D. Higher metric scores indicate better model performance. Besides, following [10], we adopt an average metric $S^{*}_{m}$ [22] as follows:
$S^{*}_{m}=\frac{1}{4}\mathrm{(BLEU\text{-}4+ROUGE_{L}+METEOR+CIDEr\text{-}D)}$ (16)
III-B Training of RSCaMa
III-B1 Objective Function
We adopt a supervised learning way to train our model. The cross-entropy function is used to compute the loss between predicted captions and ground truth captions. The formulation of the captioning loss can be expressed as follows:
$\mathcal{L}_{\text{cap}}=-\frac{1}{N}\sum_{n=1,v=1}^{N,V}y_{n}^{(v)}\log(p_{n}^{(v)})$ (17)
Here, $N$ represents the total number of word tokens, $V$ denotes the vocabulary size, $y_{n}$ denotes the one-hot vector representation of the $n$-th word in the reference sentence, and $p_{n}$ is probability vector of the predicted word.
III-B2 Implementation Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we use the Adam optimizer to optimize the model parameters. The dimension of the word vector is 768. The number of CaMa layers is set to 3. Our final RSCaMa uses one Transformer’s decoder layer with the cross-attention as the language decoder because of its better performance.
III-C Comparison to State-of-the-Art
Currently, mainstream methods commonly utilize CNNs and Transformers as basic deep learning components of the “neck” design to efficiently process bi-temporal features extracted by the backbone. Methods like Capt-Rep-Diff [6], Capt-Att [6], Capt-Dual-Att [6], and DUDA [6] combine CNNs and attention mechanisms to process bi-temporal features and employ LSTM for language decoding. MCCFormer-S [7], MCCFormer-D [7], RSICCFormer [5], PSNet [9] and PromptCC [10] use Transformer as the core component to construct the neck and employ Transformer’s decoder to generate captions. In this section, we compare RSCaMa with these state-of-the-art (SOTA) methods.
As shown in Table I, RSCaMa demonstrates outstanding performance, particularly excelling in key metrics such as BLEU-4 and $S^{*}_{m}$. BLEU-4 measures the similarity of 4-grams between predicted and reference sentences. Specifically, compared to the latest PromptCC, our model shows +1.70% on BLEU-4 and +1.11% on $S^{*}_{m}$. This significant improvement enhancement is attributed to our innovative spatial-temporal modelling units: SD-SSM and TT-SSM. Their collaborative effect enhances the model’s capability to capture spatial-temporal dynamics. Experimental results also suggest that Mamba, as an emerging deep-learning model, has the potential to become a significant option in future model designs. Further exploration of combinations of Mamba with CNNs, Transformers, and Multilayer Perceptrons (MLPs) will be a promising research direction to advance RSICC technology.
TABLE I: Comparisons experiments on the LEVIR-CC dataset, where the bolded results are the best. Different from previous methods relying on CNNs or Transformers as basic elements of the model, combined with various attention mechanisms, our RSCaMa introduce mamba into RSICC for the first time.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
Capt-Rep-Diff [6] 72.90 61.98 53.62 47.41 34.47 65.64 110.57 64.52
Capt-Att [6] 77.64 67.40 59.24 53.15 36.58 69.73 121.22 70.17
Capt-Dual-Att [6] 79.51 70.57 63.23 57.46 36.56 70.69 124.42 72.28
DUDA [6] 81.44 72.22 64.24 57.79 37.15 71.04 124.32 72.58
MCCFormer-S [7] 79.90 70.26 62.68 56.68 36.17 69.46 120.39 70.68
MCCFormer-D [7] 80.42 70.87 62.86 56.38 37.29 70.32 124.44 72.11
RSICCFormer [5] 84.72 76.27 68.87 62.77 39.61 74.12 134.12 77.65
PSNet [9] 83.86 75.13 67.89 62.11 38.80 73.60 132.62 76.78
PromptCC [10] 83.66 75.73 69.10 63.54 38.82 73.72 136.44 78.13
RSCaMa 85.79 77.99 71.04 65.24 39.91 75.24 136.56 79.24
TABLE II: Ablation studies on the SD-SSM and TT-SSM, where the bolded results are the best.
Method SD-SSM TT-SSM BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
Baseline 85.14 77.23 70.37 64.75 39.40 74.38 135.28 78.45
85.75 77.85 70.85 65.01 39.67 74.90 135.91 78.87
RSCaMa 85.79 77.99 71.04 65.24 39.91 75.24 136.56 79.24
TABLE III: Ablation studies on three language decoders. Our RSCaMa performs best when using the Transformer decoder with cross-attention.
Method BLEU-4 METEOR ROUGEL CIDEr-D $S^{*}_{m}$
Mamba 63.13 37.83 72.61 127.26 75.21
GPT-style 64.04 39.36 74.10 134.64 78.04
Trans-Dec 65.24 39.91 75.24 136.56 79.24
III-D Ablation Studies
III-D1 CaMa Layers
The CaMa layers consist of SD-SSM and TT-SSM. Table II validates the effectiveness of these two modules. The structure of our baseline model is similar to SD-SSM, but the key difference lies in using self-features rather than differential features to multiply with the output of bidirectional SSM, following the original design of Mamba. Experimental results indicate that SD-SSM and TT-SSM significantly contribute to improving change captioning performance. This is because SD-SSM effectively enhances the model’s understanding of spatial changes by multiplying differential features with the output of bidirectional SSM, while TT-SSM promotes effective temporal interaction of bi-temporal features through a token-wise cross-scanning manner.
III-D2 Language Decoder
RSICC require transforming visual change into language. The language decoders adopted by previous methods have evolved with the development of language models, from recurrent neural networks to Transformers. As a novel language model, the performance of Mamba in cross-modal understanding is yet to be explored. Table III compares three different language decoding methods: the Mamba language model, the Transformer decoder with cross-attention, and the GPT-style decoder with causal attention only. In our experiments, for the first two methods, we prefix the visual embeddings to the language decoders to generate words one by one. For the last method, visual embeddings are integrated into the decoding process as keys (K) and values (V) of the cross-attention mechanism. The experimental results in Table III demonstrate that the Transformer decoder with cross-attention mechanism performs best in transforming visual to textual information, surpassing the simple visual embedding prefix strategy (i.e., Mamba and GPT-style decoder), thus proving the superiority of cross-attention in cross-modal understanding. In the future, it is promising to explore the combination of Mamba with novel attention mechanisms rather than the simple visual prefix strategy to enhance its cross-modal understanding ability when generating each word.
Figure 4: Captioning results on the LEVIR-CC dataset. Sentence (a) is one of the five ground-truth sentences. Sentence (b) is from the baseline, while (c) is from our RSCaMa. More accurate and detailed words are marked in green. Red words are not accurate.
III-E Qualitative Results
Fig. 4 showcases some captioning results. Each image pair is accompanied by one of the five ground-truth sentences, along with the sentences generated by the baseline model and our RSCaMa. Notably, our RSCaMa demonstrates a remarkable ability to accurately distinguish changes in the scenes, as illustrated by the first image pair. Furthermore, our RSCaMa excels in identifying the categories of changing objects, their locations, and the relationships between objects. For instance, in the second image pair, our RSCaMa accurately identifies the number of the changed house, while in the third image pair, it correctly identifies the changed building’s location in the lower right corner.
IV Conclusion
In this paper, we introduce Mamba (a state space model) into the RSICC task and propose a novel approach called RSCaMa. RSCaMa iteratively processes bi-temporal features from spatial and temporal domains using multiple CaMa layers composed of SD-SSM and TT-SSM. SD-SSM uses differential features to enhance change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner. Experimental results demonstrate the superiority of RSCaMa over existing methods, highlighting the potential of Mamba in RSICC tasks. Besides, our systematic evaluation of three language decoding schemes provides valuable insights for future research.